# The PLATO Mission

### v1

@@ Abstract
PLATO (PLAnetary Transits and Oscillations of stars) is ESA’s M3 mission designed to detect and characterise extrasolar planets and perform asteroseismic monitoring of a large number of stars. PLATO will detect small planets (down to 2 R) around bright stars (11 mag), including terrestrial planets in the habitable zone of solar-like stars. With the complement of radial velocity observations from the ground, planets will be characterised for their radius, mass, and age with high accuracy (5%, 10%, 10% for an Earth-Sun combination respectively). PLATO will provide us with a large-scale catalogue of well-characterised small planets up to intermediate orbital periods, relevant for a meaningful comparison to planet formation theories and to better understand planet evolution. It will make possible comparative exoplanetology to place our Solar System planets in a broader context. In parallel, PLATO will study (host) stars using asteroseismology, allowing us to determine the stellar properties with high accuracy, substantially enhancing our knowledge of stellar structure and evolution.
The payload instrument consists of 26 cameras with 12cm aperture each. For at least four years, the mission will perform high-precision photometric measurements. Here we review the science objectives, present PLATO´s target samples and fields, provide an overview of expected core science performance as well as a description of the instrument and the mission profile at the beginning of the serial production of the flight cameras. PLATO is scheduled for a launch date end 2026. This overview therefore provides a summary of the mission to the community in preparation of the upcoming operational phases.
@@ 4 Stellar Science
- 5 The PLATO Complementary Science programs
@@ 1 Introduction
The PLATO mission (PLAnetary Transits and Oscillations of stars) is able to detect and characterise a large number of exoplanetary systems, including terrestrial exoplanets orbiting bright solar-type stars in their habitable zone. PLATO was selected as ESA’s M3 mission in the Cosmic Vision 2015-2025 Programme in 2014 [254]. PLATO satellite data will provide accurate and precise planetary radii and architectures of a large number of planetary systems via the photometric transit method. The light curve data will also enable accurate stellar parameters, including evolutionary ages, to be derived via asteroseismic analysis. In combination with high-precision spectroscopic ground-based follow-up observations, accurate planetary masses, and hence mean planetary densities, will be determined.
@@
The PLATO satellite consists of an ESA-provided satellite platform and a payload including 26 cameras, on-board computers and power supply units jointly developed by ESA and the international PLATO Mission Consortium (PMC). Launch is foreseen for end 2026. The cameras will operate in white light and obtain high-precision photometric light curves of thousands of stars.
The initial PLATO mission layout and its science objectives are described in [254], presenting the status just before mission selection. In the present paper we summarise the status of the consolidated mission design at the beginning of flight model production and show the mission in the context of current exoplanetary research. We provide an overview of the expected impact in exoplanet science including planet yields and performance for planet characterisation (Section 3). Section 4 discusses PLATO’s impact on stellar science, including a discussion on accuracy versus precision and the prospects for detecting solar-like oscillations in PLATO samples. The potential of PLATO concerning complementary sciences beyond its core programme is discussed in Section 5. Section 6 defines the PLATO P1 to P5 stellar samples and presents options for the distribution of on-board processed light curves versus imagettes for detailed analysis on ground. The status of target field and observing mode selection is summarised in Section 7. The science ground segment, PLATO data products and data releases are presented in Section 8. The organised ground-based follow-up to determine masses for planets in PLATO’s prime sample is a key element of the mission (Section 9). In addition to its core science goals, PLATO will provide a wealth of complementary science data available via the Guest Observer programme, which will be part of the legacy of the mission, see Section 10. The expected instrument signal-to-noise levels as well as the mission and payload designs are highlighted in Sections 11, 12 and 13. Finally, the synergies with other missions are presented in 14.
@@ 2 Overview of PLATO Science Objectives
Among the key goals of PLATO are high accuracy parameters for planets orbiting F5-K7 dwarf and sub-giant stars. Note that we use the term “accurate” planet parameters for those cases where stellar parameters can be derived with highest precision and stellar models are well constrained (see Section 4). For an Earth-sized planet (Earth radius: re) orbiting a 10 mag G0V star an accuracy of 3% in planetary radius (5% for a 11 mag star) and 10% in stellar age shall be derived. This configuration is used as the benchmark for the mission. Better precision is expected for larger planets or brighter host stars (see Section 3.2). In addition, PLATO will observe stars that are sufficiently bright to enable the determination of planetary masses (Earth mass: me) with an accuracy of 10% for a terrestrial planet (defined as 2 Re and 10 Me) orbiting a G0V star via the radial velocity method from ground-based follow-up observations. The resulting planetary bulk properties allow us to explore the diversity of planets and identify typical planet populations, which in turn help better constrain planet formation models. Small planets with high mean densities in the HZ will be prime candidates for potentially habitable planets and allow an estimation of , here defined as the fraction of Earth-like planets in the HZ per host star. Furthermore, precise measurements for close-in high-density planets will show how many have a Mercury-like interior structure which will allow a study of their refractory (e.g. Fe and Si) content and help test formation theories for this class of planets. Extending our knowledge towards longer orbital periods will show whether extreme-density small planets also exist on such orbits. Other questions which can be addressed with PLATO data (and combined with e.g. stellar properties) include, “What is a typical internal structure and composition of terrestrial and mini-gas planets? How does it evolve due to stellar interactions (losses)? What is the core size of gaseous planets?”.
@@
The age of planetary systems is a key parameter to explore how planetary system properties evolve and provides observational constraints to formation models as well as to how gravitational instabilities deplete planetary systems over time [308]. Furthermore, details of the evolutionary stages of planets are only known from one example so far, our Solar System. In particular for terrestrial planets and their complex evolution as well as for the new and poorly known class of mini-Neptunes better constraints can be provided once a sample of planets observed at different evolution stages becomes available. Open questions are, e.g., “When does the magma ocean phase for terrestrial planets end?” and “Are our model timescales for the evolution of gaseous planets correct?”. To further our understanding of these issues, homogeneous and precise age determination is required, an issue we are still facing today with the various age indicators. The seismic characterisation of a large sample of bright stars across the Hertzsprung-Russell (HR) diagram will lead to significantly improved stellar models, allowing for substantially more reliable age characterisation of stars in general.
@@
The architecture of planetary systems includes parameters such as the distribution of planet masses and types (terrestrial or gaseous) over orbital separation, the co-planarity of systems, and orbital parameters such as, e.g., orbital eccentricities (e.g. Van Eylen and Albrecht 301, Van Eylen et al 303) and inclinations. Key questions to address include, e.g., “What fraction of planetary systems have a structure similar to the Solar System? How many have multiple/no gas giants? Was migration caused by planet-planet interaction or disk migration?”. First analyses of Kepler mission results, but also comparisons with planet synthesis models (e.g. Borucki et al 47, Lissauer et al 195, Weiss et al 316, Millholland et al 216, Wang and Ji 315, Mishra et al 217) suggest adjacent planets having similar masses and sizes with a trend towards increasing masses for outer planets. Dynamical interactions between planet embryos can lead to typical spacing between planets with small planets being more densly packed than large/massive ones. Today, these apparent trends are based on only a few tens of well-characterised planetary systems in a limited parameter space [e.g., 217] and there are many exceptions to the trends. Furthermore, the dynamical evolution and development of, e.g., orbital resonances during early ages of planetary systems need to be better understood and linked to system properties. Nevertheless, the available early studies already show the large potential for well-characterised samples of planet system architectures to constrain planet formation theories. PLATO will significantly extend the number of well-characterised planetary systems, in particular for longer orbital periods longer than 80 days. These longer periods are underrepresented in past and ongoing mission data (see Section 3.1.2 for a comparison to the Transiting Exoplanet Survey Satellite (TESS, Ricker et al 259).
@@
Occurrence rates for different types of planets around different types of stars have been provided by both transit and radial velocity surveys (e.g. Howard et al 151, Fressin et al 114, Christiansen et al 77, Deleuil et al 99, Kunimoto et al 183, see Section 3.1.2 for more). A proper estimate of an occurrence rate requires however a good estimate of some key parameters such as the survey detection sensitivity to estimate its completeness, the planet parameters, and a good knowledge of the underlying stellar population. The latter is a serious limitation that prevents direct comparisons of published occurrence rates and limits the reliability of the estimates. PLATO will benefit from the in-depth characterisation of the stellar fields the instrument will observe and an analysis with the most up-to-date stellar models. This will allow for the exploration of possible correlations of planet parameters with stellar properties. For example, the chemical element ratios of a stellar photosphere is thought to be a good first proxy of its planets’ initial compositions alleviating the degeneracies that occur when deriving the planet’s composition from its radius and mass only [101, 52, 1]. This hypothesis has been recently questioned by various studies [251, 276, 4] which highlight discrepancies between the actual planetary composition and what is expected from a primordial origin as reflected by the chemical ratios of the stars. In addition, stellar properties such as activity are key parameters for our understanding of potential planet habitability. In particular, stellar rotational evolution is a key ingredient to understand planetary migration through tidal and magnetic interaction (e.g. Lai 184, Mathis et al 203, Strugarek et al 289, 290, Damiani and Mathis 95, Ahuir et al 8).
@@
Determining planet host star parameters requires improving today’s stellar models and stellar evolution theory in general. PLATO light curve data will be used to measure the oscillation frequencies of stars, which will be interpreted via asteroseismic modelling [e.g., 119, 5, for recent reviews] to test evolution theory. Stellar models constrained by asteroseismology and cross-calibration with classical methods of stellar modelling will be key in obtaining planet host star parameters to address the overall PLATO science goals.
In the absence of seismic age determination, measurements of rotation periods from the PLATO light curves will allow us to estimate the age of a given target through gyrochronology [284, 25, 26] for stars like solar-type stars which are losing angular momentum through the stellar wind due to their magnetic activity. The PLATO rotation period catalogue will complete and extend the catalogues already available for Kepler/K2/TESS (e.g. Nielsen et al 231, McQuillan et al 210, Santos et al 270, 271, Holcomb et al 150). It is also expected to detect and characterise activity cycles which are shorter than the solar cycle, such as the Rieger-like cycle [260, 136]. For stars with planets it is furthermore important to understand how activity and rotation is affected by close-in planets, a task for which PLATO will expand our currently available data base.
@@
Planets identified around the brightest stars will likely be “Rosetta Stones” for spectroscopic follow-up to study their atmospheric structure and composition. Apart from the gaseous planets, those low-mass PLATO planets accessible to e.g. JWST (James-Web-Space-Telescope), or its successors, have the potential to provide new insights into the link between mantle composition, iron-to-silicate ratio, redox state and atmospheric compositions. They would allow, e.g., the identification of secondary atmospheres (water, CO2) on such low-mass planets (e.g., Ortenzi et al 239, Kabath et al 168), which are barely distinguishable from just mass and radius modelling alone.
@@
The following chapters detail PLATO’s science objectives and provide analysis of the respective accuracy and precision that the instrument can achieve.
@@ 3.1.1 State of the Art
The pioneering early discoveries made by ground-based telescopes (e.g. WASP [252], HATNet [22]) and the CoRoT space mission [21, 99] provided a first glimpse on the diversity of extrasolar giant and close-in planet properties. It was, however, the Kepler mission [179, 48] which enlarged the sample of small-sized planets and enabled the first studies on exoplanet population properties [28]. For ensemble properties to become apparent, a sufficiently large number of planets with well-known parameters are needed. Today, TESS data provide a wealth of information on the population of planets with orbital periods up to about 100 days [135].
Analyses of the planets with the most accurately determined radii (i.e. those with the best characterised stars) reveal a non-homogeneous distribution in bulk planet properties. Whereas in our Solar System the distribution of planetary radii follow a power law, Kepler data show that there is structure in the distribution of exoplanet systems when evaluating the planet population at large. This is particularly interesting for small planets. While there are still questions over completeness at small radii, a deficit of planets at 1.8 (Fig. 1) suggests physical processes (photo evaporation or core-powered mass loss) that separate the super-earth and sub-neptune populations - the so-called radius valley [199, 116, 115, 307, 161, 149]. The benefit of well-characterised host stars from asteroseismology when investigating the slope of the radius valley with increasing orbital distance was shown by [302]. However, although our current knowledge on exoplanet populations already provides exciting results challenging theories of planetary system formation and evolution, it is still heavily affected by observational biases. For example, the population of small planets represented in Figure 1 is dominated by planets with orbital periods of less than 80-100 days. There is a clear need to explore whether and how the radius valley extends to larger orbits. PLATO data will be excellent to analyze such trends at intermediate orbital distances.
@@
Figure 2 illustrates the anticipated detection range of PLATO with respect to the location of the HZ of solar-like and M dwarf host stars. Only planets with measured radii and masses are shown in the figure. A few characterised small planets in the HZ of cool M dwarf host stars are already known (e.g. the TRAPPIST-1 system (Gillon et al 127) and more detections around cool stars are expected in the near future from NASA’s TESS mission. In addition, ESA’s CHEOPS (CHaracterising ExOPlanet Satellite) mission [39] can provide higher-precision radii for known exoplanets and therefore improve our knowledge of well-characterised bodies. However, these missions cover orbital distance ranges well below 1 AU.
@@
Our current best knowledge on planet occurrence rates results from homogeneous re-analyses of Kepler/K2 data [182, 53]. Derived rates (number of planets per star) are in the range of a few to less than one percent, depending on size and orbital period. However, for small, long-period planets only upper limits can be given and uncertainties are large. Table 1 provides an overview of published values on the average number of Earth-like planets in the HZ. Uncertainties are large, whether they address solar-like or M dwarf host stars. For example, Kunimoto and Matthews [182] provide upper limits from Kepler data reaching from few percent up to 40% for Earth-analogues (see their Figures 4 to 7). It is one of the prime goals of PLATO to derive an improved estimate of the occurrence rate of well-characterised small planets in the HZ of solar-like stars.

### v2

@@ Abstract
PLATO (PLAnetary Transits and Oscillations of stars) is ESA’s M3 mission designed to detect and characterise extrasolar planets and perform asteroseismic monitoring of a large number of stars. PLATO will detect small planets (down to 2REarth) around bright stars (11 mag), including terrestrial planets in the habitable zone of solar-like stars. With the complement of radial velocity observations from the ground, planets will be characterised for their radius, mass, and age with high accuracy (5%, 10%, 10% for an Earth-Sun combination respectively). PLATO will provide us with a large-scale catalogue of well-characterised small planets up to intermediate orbital periods, relevant for a meaningful comparison to planet formation theories and to better understand planet evolution. It will make possible comparative exoplanetology to place our Solar System planets in a broader context. In parallel, PLATO will study (host) stars using asteroseismology, allowing us to determine the stellar properties with high accuracy, substantially enhancing our knowledge of stellar structure and evolution.
The payload instrument consists of 26 cameras with 12cm aperture each. For at least four years, the mission will perform high-precision photometric measurements. Here we review the science objectives, present PLATO´s target samples and fields, provide an overview of expected core science performance as well as a description of the instrument and the mission profile towards the end of the serial production of the flight cameras. PLATO is scheduled for a launch date end 2026. This overview therefore provides a summary of the mission to the community in preparation of the upcoming operational phases.
@@ 4 Stellar Science
- 5 The PLATO Complementary Science
@@ 1 Introduction
The PLATO mission (PLAnetary Transits and Oscillations of stars) is designed to detect and characterise a large number of exoplanetary systems, including terrestrial exoplanets orbiting bright solar-type stars in their habitable zone. PLATO was selected as ESA’s M3 mission in the Cosmic Vision 2015-2025 Programme in 2014 [278]. PLATO satellite data will provide accurate and precise planetary radii and architectures of a large number of planetary systems via the photometric transit method. The light curve data will also enable accurate stellar parameters, including evolutionary ages, to be derived via asteroseismic analysis. In combination with high-precision spectroscopic ground-based follow-up observations, accurate planetary masses, and hence mean planetary densities, will be determined.
@@
The PLATO satellite consists of an ESA-provided satellite platform and a payload including 26 cameras (see Section 13). Of these, 24 cameras will operate in white light and obtain high-precision photometric light curves of thousands of stars, and 2 cameras will provide colour information by observing in the blue (505-700 nm) and in the red (665-1000 nm) wavelength range respectively. Adding to this are the on-board computers and power supply units. The PLATO payload is developed by the international PLATO Mission Consortium (PMC) including contributions from ESA. Launch is foreseen for end 2026.
The initial PLATO mission layout and its science objectives are described in [278], presenting the status just before mission selection. In the present paper we summarise the status of the consolidated mission design towards the end of flight model production and show the mission in the context of current exoplanetary research.
However, let us first have a look at the major updates on the mission since 2014. The scientific goals and research questions described in [278], and which can be addressed with PLATO, are still valid today. After the successful launch of NASA´s TESS mission in 2018, PLATO puts, however, less emphasize on short-period planets and short pointing target fields, although the option to do so is kept in the mission requirements. The focus of PLATO´s nominal mission is clearly on long orbital period planets now. The major design change since 2014 concerns the move from 32 "Normal"-cameras to 24 N-Cams in the current design. See Section 7 for the expected science performances with the current design. There have been multiple further updates, refinements and adaptations of the payload design made since 2014, which are however too detailed to be listed here. The overall design of the cameras with their support units remained, however, similar to the concept presented in 2014. Concerning the mission itself, the final satellite design and choice of industry prime of ESA was made from three competing studies during the Phase B study phase.
In this review we provide an overview of the expected impact of PLATO in exoplanet science including planet yields and performance for planet characterisation (Section 3). Section 4 discusses PLATO’s impact on stellar science, including a discussion on accuracy versus precision and the prospects for detecting solar-like oscillations in PLATO samples. The potential of PLATO concerning complementary science topics beyond its core programme is discussed in Section 5. Section 6 defines the PLATO P1 to P5 stellar samples and presents options for the distribution of on-board processed light curves versus imagettes for detailed analysis on ground. The status of target field and observing mode selection is summarised in Section 7. The science ground segment, PLATO data products and data releases are presented in Section 8. The organised ground-based follow-up to determine masses for planets in PLATO’s prime sample is a key element of the mission (Section 9). In addition to its core science goals, PLATO will provide a wealth of complementary science from its observed samples P1 to P5 and its science calibration stars. Moreover, PLATO will also offer an extensive Guest Observer programme, which will be part of the legacy of the mission, see Section 10. The expected instrument signal-to-noise levels as well as the mission and payload designs are highlighted in Sections 11, 12 and 13. Finally, the synergies with other missions are presented in 14.
@@ 2 Overview of PLATO Science Objectives
Among the key goals of PLATO are high accuracy parameters for planets orbiting F5-K7 dwarf and sub-giant stars. Note that we use the term “accurate” planet parameters for those cases where stellar parameters can be derived with highest precision and stellar models are well constrained (see Section 4). For an Earth-sized planet (Earth radius: REarth) orbiting a 10 mag G0V star an accuracy of 3% in planetary radius (5% for a 11 mag star) and 10% in stellar age shall be derived (see also Table 5). This configuration (for planets up to 1 AU) is used as the benchmark for the mission. Better performance is expected for larger planets or brighter host stars (see Section 3.2). In addition, PLATO will observe stars that are sufficiently bright to enable the determination of planetary masses (Earth mass: MEarth) with an accuracy of 10% for a terrestrial planet (defined as 2 REarth and 10 MEarth) orbiting a G0V star via the radial velocity method from ground-based follow-up observations. The resulting planetary bulk properties allow us to explore the diversity of planets and identify typical planet populations, which in turn help better constrain planet formation models. Small planets with high mean densities in the HZ will be prime candidates for potentially habitable planets and allow an estimation of , here defined as the fraction of Earth-like planets in the HZ per host star. Furthermore, precise measurements for close-in high-density planets will show how many have a Mercury-like interior structure which will allow a study of their refractory (e.g. Fe and Si) content and help test formation theories for this class of planets. Extending our knowledge towards longer orbital periods will show whether extreme-density small planets also exist on such orbits. Other questions which can be addressed with PLATO data (and combined with e.g. stellar properties) include, “What is a typical internal structure and composition of terrestrial and mini-gas planets? How does it evolve due to stellar interactions (losses)? What is the core size of gaseous planets?”.
@@
The age of planetary systems is a key parameter to explore how planetary system properties evolve and provides observational constraints to formation models as well as to how gravitational instabilities deplete planetary systems over time [332]. Furthermore, details of the evolutionary stages of planets are only known from one example so far, our Solar System. In particular for terrestrial planets and their complex evolution as well as for the new and poorly known class of mini-Neptunes better constraints can be provided once a sample of planets observed at different evolution stages becomes available. Open questions are, e.g., “When does the magma ocean phase for terrestrial planets end?” and “Are our model timescales for the evolution of gaseous planets correct?”. To further our understanding of these issues, homogeneous and precise age determination is required, an issue we are still facing today with the various age indicators. The seismic characterisation of a large sample of bright stars across the Hertzsprung-Russell (HR) diagram will lead to significantly improved stellar models, allowing for substantially more reliable age characterisation of stars in general. Gyrochronology, magneto-gyrochronology [228] and age-activity relationships (see Sect. 4.6) will be used to provide estimates for the ages of targets where p-mode oscillations cannot be detected, notably those beyond spectral type K2-K3. This will considerably increase the statistical samples upon which the evolution of planetary systems can be investigated.
@@
The architecture of planetary systems includes parameters such as the distribution of planet masses and types (terrestrial or gaseous) over orbital separation, the co-planarity of systems, and orbital parameters such as, e.g., orbital eccentricities (e.g. Van Eylen and Albrecht 325, Van Eylen et al 327) and inclinations. Key questions to address include, e.g., “What fraction of planetary systems have a structure similar to the Solar System? How many have multiple/no gas giants? Was migration caused by planet-planet interaction or disk migration?”. First analyses of Kepler mission results, but also comparisons with planet synthesis models (e.g. Borucki et al 49, Lissauer et al 214, Weiss et al 340, Millholland et al 238, Wang and Ji 339, Mishra et al 239) suggest adjacent planets having similar masses and sizes with a trend towards increasing masses for outer planets [e.g., 84]. Dynamical interactions between planet embryos can lead to typical spacing between planets with small planets being more densely packed than large/massive ones. Today, these apparent trends are based on only a few tens of well-characterised planetary systems in a limited parameter space [e.g., 239] and there are many exceptions to the trends. Furthermore, the dynamical evolution and development of, e.g., orbital resonances during early ages of planetary systems need to be better understood and linked to system properties. Nevertheless, the available early studies already show the large potential for well-characterised samples of planet system architectures to constrain planet formation theories. PLATO will significantly extend the number of well-characterised planetary systems, in particular for orbital periods longer than 80 days. These longer periods are underrepresented in past and ongoing mission data (see Section 3.1.2 for a comparison to the Transiting Exoplanet Survey Satellite (TESS, Ricker et al 282).
@@
Occurrence rates for different types of planets around different types of stars have been provided by both transit and radial velocity surveys (e.g. Howard et al 165, Fressin et al 125, Christiansen et al 82, Deleuil et al 104, Kunimoto et al 198, see Section 3.1.2 for more). A proper estimate of an occurrence rate requires however good knowledge of some key parameters such as the survey detection sensitivity to estimate its completeness, the planet parameters, and a good overview of the underlying stellar population. The latter is a serious limitation that prevents direct comparisons of published occurrence rates and limits the reliability of the estimates. PLATO will benefit from the in-depth characterisation of the stellar fields the instrument will observe as well as an analysis with the most up-to-date stellar models. This will allow for the exploration of possible correlations of planet parameters with stellar properties. For example, the chemical element ratios of a stellar photosphere are thought to be a good first proxy of its planets’ initial compositions alleviating the degeneracies that occur when deriving the planet’s composition from its radius and mass only [109, 57, 1]. This hypothesis has been recently questioned by various studies [275, 300, 4] which highlight discrepancies between the actual planetary composition and what is expected from a primordial origin as reflected by the chemical ratios of the stars. In addition, stellar properties such as activity are key parameters for our understanding of potential planet habitability because it impacts, among other, on the evaporation and the chemical evolution of planetary atmospheres. In particular and as discussed in Sect.4.6, stellar rotational evolution is a key ingredient to understand planetary migration through tidal and magnetic interaction (e.g. Lai 200, Mathis et al 224, Strugarek et al 311, 312, Damiani and Mathis 100, Ahuir et al 9).
@@
Determining planet host star parameters requires improving today’s stellar models and stellar evolution theory in general. PLATO light curve data will be used to measure the oscillation frequencies of stars, which will be interpreted via asteroseismic modelling [e.g., 130, 5, for recent reviews] to test evolution theory. Stellar models constrained by asteroseismology and cross-calibration with classical methods of stellar modelling will be key in obtaining planet host star parameters to address the overall PLATO science goals.
In the absence of seismic age determination, measurements of rotation periods from the PLATO light curves will allow us to estimate the age of a given target through gyrochronology [307, 24, 25] for stars like solar-type stars which are losing angular momentum through the stellar wind due to their magnetic activity. The PLATO rotation period catalogue will complete and extend the catalogues already available for Kepler/K2/TESS (e.g. Nielsen et al 254, McQuillan et al 232, Santos et al 293, 294, Holcomb et al 164). It is also expected to detect and characterise activity cycles which are shorter than the solar cycle, such as the Rieger-like cycle [283, 149]. For stars with planets it is furthermore important to understand how activity and rotation is affected by close-in planets, a task for which PLATO will expand our currently available data base.
@@
Planets identified around the brightest stars will likely be “Rosetta Stones” for spectroscopic follow-up to study their atmospheric structure and composition. Apart from the gaseous planets, those low-mass PLATO planets accessible to e.g. JWST (James-Web-Space-Telescope), or its successors, have the potential to provide new insights into the link between mantle composition, iron-to-silicate ratio, redox state and atmospheric compositions. They would allow, e.g., the identification of secondary atmospheres (water, CO2) on such low-mass planets (e.g., Ortenzi et al 259, Kabath et al 179), which are barely distinguishable from just mass and radius modelling alone.
@@
The following sections detail PLATO’s science objectives and provide analysis of the respective accuracy and precision that the instrument can achieve.
@@ 3.1.1 State of the Art
The pioneering early discoveries made by ground-based telescopes (e.g. WASP, Pollacco et al 276; HATNet, Bakos et al 23) and the CoRoT space mission [22, 104] provided a first glimpse on the diversity of extrasolar giant and close-in planet properties. It was, however, the Kepler mission [192, 50] which enlarged the sample of small-sized planets and enabled the first studies on exoplanet population properties [28]. For ensemble properties to become apparent, a sufficiently large number of planets with well-known parameters are needed. Today, TESS data provide a wealth of information on the population of planets with orbital periods up to about 100 days [147].
Analyses of the planets with the most accurately determined radii (i.e. those with the best characterised stars) reveal a non-homogeneous distribution in bulk planet properties. Whereas in our Solar System the distribution of planetary radii follow a power law, Kepler data show that there is structure in the distribution of exoplanet systems when evaluating the planet population at large. This is particularly interesting for small planets. While there are still questions over completeness at small radii, a deficit of planets at 1.8 REarth (Fig. 1) suggests physical processes (photo evaporation or core-powered mass loss) that separate the super-earth and sub-neptune populations - the so-called radius valley [219, 127, 263, 126, 331, 173, 163]. The benefit of well-characterised host stars from asteroseismology when investigating the slope of the radius valley with increasing orbital distance was shown by [326]. However, although our current knowledge on exoplanet populations already provides exciting results challenging theories of planetary system formation and evolution, it is still heavily affected by observational biases. For example, the population of small planets represented in Figure 1 is dominated by planets with orbital periods of less than 80-100 days. There is a clear need to explore whether and how the radius valley extends to larger orbits. PLATO data will be excellent to analyze such trends at intermediate orbital distances.
@@
Figure 2 illustrates the anticipated detection range of PLATO with respect to the location of the HZ of solar-like and M dwarf host stars. Only planets with measured radii and masses are shown in the figure. A few characterised small planets in the HZ of cool M dwarf host stars are already known (e.g. the TRAPPIST-1 system Gillon et al 138, but see also LHS-1140, Dittmann et al 107, Lillo-Box et al 213, Cadieux et al 66) and more detections around cool stars are expected in the near future from NASA’s TESS mission, including planets in the habitable zone, [181, 196]. In addition, ESA’s CHEOPS (CHaracterising ExOPlanet Satellite) mission [39] can provide higher-precision radii for known exoplanets and therefore improve our knowledge of well-characterised bodies. However, these missions cover orbital distance ranges well below 1 AU.
@@
Our current best knowledge on planet occurrence rates results from homogeneous re-analyses of Kepler/K2 data [197, 58]. Derived rates (number of planets per star) are in the range of a few percent down to less than one percent, depending on size and orbital period. However, for small, long-period planets only upper limits can be given and uncertainties are large. Table 1 provides an overview of published values on the average number of Earth-like planets in the HZ. Uncertainties are large, whether they address solar-like or M dwarf host stars. For example, Kunimoto and Matthews [197] provide upper limits from Kepler data reaching from few percent up to 40% for Earth-analogues (see their Figures 4 to 7). It is one of the prime goals of PLATO to derive an improved estimate of the occurrence rate of well-characterised small planets in the HZ of solar-like stars.